sin x=1/4 and 0 < (or equal to) x < (or equal to) pie/2, the cos x=?
1. square root 7/ 4
2. Square root 15/ 4
3. 3/4
4. 1/ square root of 15
5. 1/ square root of 17

Answer by fractalier(6550)

We know that sin^2(x) + cos^2(x) = 1, so that
cos^2(x) = 1 - sin^2(x)
cos^2(x) = 1 - (1/4)^2 = 1 - 1/16
cos^2(x) = 15/16
cos(x) = sqrt(15) / 4
Choice 2.